According to the ECC website, the average age of ECC students is 28.2 years. Suppose we collect data from 30 Mth120 online students and find an average age of 20.3 years, with a standard deviation of 2.3 years.

Is there enough evidence at the 5% level of significance to support the claim that the average age of Mth120 online students is less than 28.2 years?
Answer #1

The Hypotheses are:

Rejection region at 0.05 level of significance :

Reject Ho if T< -t 0.05,29 =-1.70

Test Statistic:

Since the sample standard deviation is given hence we will be using T- statistic to check the Hypothesis as

Conclusion:

Since The test statistic T <<-t =-1.65 hence it falls in rejection region so, we reject the null Hypothesis and conclude that there is enough evidence to support the claim at 0.05 level of significance.

T-Table for t value calculation at Degree of freedom=n-1=30-1=29